# Disentangling Boltzmann brains, the time-asymmetry of memory, the H theorem, and the second law

### v1

@@ IV Discussion
In this paper, we have elaborated on the following contributions:
@@
    The standard argument in favor of the BB hypothesis contradicts itself by implicitly assuming the second law;
@@
    Boltzmann showed that the H function evolves as a Boltzmann process, as defined above.
@@
    Boltzmann did not prove that the H function evolves as some specific singleton-equivalent stochastic process.
@@
    Reasoning based on the correct understanding of the H theorem has nothing to say about the BB hypothesis.
@@
    We introduced formal definitions of the BB hypothesis, the PH, and a variant of the BB hypothesis that does not suffer from any form of cognitive instability, which we call the 1000CE hypothesis.
Moreover, we have disentangled the BB hypothesis, the time-asymmetry of memory, the H theorem, and the second law in a few key ways. First, we have argued that the reliability of experimental data and Type-3 memory cannot be disentangled from the second law. Second, we have shown that the H theorem is independent of each the BB hypothesis, the PH, and the 1000CE hypothesis, by virtue of being an independent ingredient in those hypotheses. Third, we have shown that those hypotheses are all versions of one another with different fluctuation times, with the caveat that the BB hypothesis is seemingly more difficult to establish priors for.
We have provided the necessary formalism to study the relationships between these concepts, using stochastic process theory, for the first time. Such a formalism is necessary to avoid the implicit assumptions that often result in circular reasoning and to highlight the central role of priors. Ultimately, the only reason we have for assigning more credence to the PH or even to the second law than we assign to the BB hypothesis is our cosmological observations concerning the big bang, ${\cal{D}}_{3}$. So before the early 20th century, when we first made those observations, we should have treated the BB hypothesis as just as probable as the second law. Moreover, whatever reliability we assign to cosmological observations implying the entropy at the big bang was low, we should arguably assign at least as much to our observations concerning the current entropy of the universe. Once we do that however, then we have a two-moment nailed set — which rules out not just the BB hypothesis, but also the naive versions of the PH and the second law currently in vogue.

### v2

@@ IV Discussion
In this paper, we introduced new stochastic process theory to allow us to disentangle the second law, the reliability of human memory, the BB hypothesis, and the Past Hypothesis (PH) with full formal rigor.
To do this we first established the following, preliminary results (some of which are similar to earlier results in the literature):
@@
    The standard argument in favor of the BB hypothesis can be seen as contradicting itself, by implicitly assuming the second law;
@@
    Boltzmann’s H theorem, properly formalized, says that the H function evolves as a particular type of stochastic process, a “Boltzmann process”, and then derive some new, useful properties of Boltzmann processes;
@@
    We point out that Boltzmann’s reasoning does not prove that the $H$ function evolves as a Boltzmann process conditioned on the value of $H$ at a single time (any more than he proved that it evolves as a Boltzmann process conditioned on the value of $H$ at two times, or zero times). This disproves the common interpretation of what Boltzmann proved;
@@
    We introduced formal definitions involving Boltzmann processes of the BB hypothesis, the Past Hypothesis (PH), and a variant of the BB hypothesis that does not suffer from any form of cognitive instability, which we call the 1000CE hypothesis;
@@
    We point out that due to (iii), the H theorem has nothing to say one way or another about the legitimacy of the BB hypothesis, nor of the second law, the PH, or the 1000CE hypothesis.
We then used these preliminary results to disentangle the BB hypothesis, the time-asymmetry of memory, the H theorem, and the second law, with more formal rigor than has been done before. Specifically, we demonstrate that neither the reliability of experimental data nor our possessing Type-3 memory can be disentangled from the second law. Second, we show that the validity of the H theorem is independent of the BB hypothesis, the PH, and the 1000CE hypothesis, since by it is an independent ingredient in those hypotheses. Third, we show that those hypotheses are all variants of one another, simply differing in the times they assume for the extreme of an entropy fluctuation — with the caveat that it the BB hypothesis has a seeming self-contradiction baked in.
Our use of stochastic process theory – never before used to analyze these issues – was necessary to avoid the implicit assumptions in many arguments in the literature, that often result in circular reasoning. Stochastic process theory also (correctly) forces us to specify the prior probabilities one must use to analyze the issues we consider.
This rigor demonstrates that, ultimately, the only reason we have for assigning more credence to the PH or even to the second law than we assign to the BB hypothesis is the credence we assign to our cosmological observations concerning the big bang. This means in turn that before the early 20th century, when we first made those cosmological observations, we should have treated the BB hypothesis as just as probable as the second law. Moreover, whatever reliability we now assign to those cosmological observations implying the entropy at the big bang was low, we should arguably assign at least as much reliability to our observations concerning the current entropy of the universe. Once we do that however, then the Boltzmann process governing the evolution of the universe’s entropy should be conditioned on two moments in time (the Big Bang and the present). However, conditioned that way, the Boltzmann process of the universe would rule out not just the BB hypothesis, but also the naive versions of the PH and the second law currently in vogue.